A ray of light travels in a block of material and exits to the water (nw = 1.33). The angle of incidence in the block is 55o . Select the CORRECT value(s) of an index of refraction for the total internal reflection to occur for the ray in the block.

A. 1.3

B. 1.5

C. 1.7

D. 2.4


1
Expert's answer
2021-07-12T12:14:12-0400
"\sin\theta=\frac{n_w}{n}\\\sin 55=\frac{1.33}{n}\\n=1.62"